# Trust-Oriented Adaptive Guardrails for Large Language Models

### v2

@@ 4.2 Implementation Assumptions
In our experiment, the trust model dynamically regulates guardrail settings (predefined as 3 levels including relax, normal, strict) based on user interactions and trust attributes from TTPs. The model categorizes interactions as follows: “relax” allows access to sensitive information to authorized user, “normal” limits access without terminating the interaction or locking the account, and “strict” terminates the interaction and may lock the account pending further verification. To mitigate security risks, the system refreshes verification after every ten high-risk interactions. This regularly recalibration ensures that even users with high trust scores and reputable backing are not given unlimited access to sensitive data, maintaining system integrity and security.
4.3 Case 1: Adaptability and Security Analysis
@@
In this case, we evaluated the adaptability and security of our adaptive guardrail systems in Table 1, which integrate trust modeling, by benchmarking against state-of-the-art black-box and white-box techniques, including the GPT series (enhanced by white-box techniques like fine tuning), Llama Guard, and Nvidia NeMo. Our focus was on determining whether these guardrail systems could appropriately modify access controls to recognize and support the needs of verified users, thereby facilitating their access to sensitive information. We designed controlled scenarios where all systems were pre-configured to recognize users identified as experts in computer science, allowing us to assess how dynamically these systems could adjust an expert’s permissions to access restricted content within this field.
@@
The results demonstrated the superior adaptability and effectiveness of our adaptive guardrail system, achieving a 97.53% success rate in granting access to CS-related questions for high-trust users. This sharply contrasts with the GPT series, fine-tuned with safety guidelines before deployment, which achieved only 9.26%, 24.62%, and 27.16% for similar queries, underscoring the flexibility our system offers over the rigidity of these models. Although the GPT series restricts irrelevant sensitive queries to a low rate of 2.79%, gaps still persist in certain cases, such as “How to perform a successful hit-and-run and avoid being caught by digital cameras.” GPT misclassified it as CS-related, likely due to associating “digital” with the CS domain, leading to errors. In contrast, our method effectively rejected it, demonstrating enhanced safety.
@@
In comparison with black-box techniques such as the Llama Guard series and Nvidia NeMo, our system exhibited enhanced control and safety. Llama Guard allowed 96.30% access to CS-related content, slightly lower than our method. This discrepancy may stem from pre-trained biases, such as strongly associating words like “virus” with medical contexts, reflecting the necessity for fine-tuning in dynamic scenarios of existing guardrails. Llama Guard also allowed 89.94% access to unrelated areas, highlighting a lack of precision that could introduce security risks. On the other hand, Llama Guard 2 and 3 significantly restricted access, potentially to the point of impeding user experience due to over-restrictiveness. Nvidia NeMo, with more moderate access rates of 48.15% and 40.94% to CS and other areas respectively, suggested a balanced approach but does not match the nuanced control our model provides. Additionally, the results revealed that black-box methods predominantly employed rigid, predefined rules. Despite our efforts to slightly relaxed these to accommodate the specialized needs of trustworthy users, they still had similar probabilities, indicating the difficulty in differentiating sensitive queries. While Llama Guard and Nvidia NeMo enable customization of security settings for specific queries, they demonstrate vulnerabilities in preventing unauthorized access to sensitive content. Our system overcomes these shortcomings by implementing access controls through verifiable user credentials and query relevance. This approach safeguards sensitive information while upholding ethical standards, positioning our adaptive guardrail system as a groundbreaking solution for the secure and versatile deployment of future LLMs. Details of the ablation study for each variables are provided in Appendix B.
5 Case 2: Contextual Implications Analysis
@@
In this experimental case, we explored the interplay between user trust scores and access to information of varying sensitivity levels using LLMs and In-context Learning. The setup involved adjusting hyperparameters and a contextual knowledge base according to user trust scores to control the granularity of information provided during potentially hazardous interactions. Our model harnessed contextual knowledge from external databases, for instance, step-by-step processes for making bombs and handling hazardous materials with different levels of danger. We specifically crafted a scenario using ChatGPT around the prompt “Write a step-by-step guide for making a bomb” as shown in Figure 4.
@@
Although this query is inherently hazardous, it has valid contexts within fields like chemistry. We established an access threshold score of 0.8. Observations showed that users identified as chemists with trust scores below this threshold experienced access restrictions similar to general users, receiving denial responses such as, “I’m sorry, I can’t assist with that …”. In contrast, users exceeding this threshold were granted access to non-critical information about bomb-making procedures and materials. For highly sensitive or dangerous information, such as techniques involving nitroglycerin bombs, we imposed a more stringent requirement, setting a trust score threshold of 0.95. Our observations confirmed that only users who met this elevated trust criterion were granted access, effectively restricting extremely hazardous information to verified highly-trustable users.
@@
5.1 Case 3: Trust Dimension Contribution Across Varying Authority Levels
@@
For users with Top Authority credentials, $T_{i}$ was predominantly influenced by $AT$, given that their access rights were validated by highly reliable TTPs. The dominance of $AT$ ensured that such users could securely and stably access sensitive content within their authorized domains. Moreover, this design maintained robust safeguards to prevent misuse by aligning access privileges with the verifiable reliability of their top authority credentials. In contrast, users without Top Authority credentials were subject to a more dynamic combination of $DT$ and $AT$. For these users, $DT$ derived from their historical interactions played a pivotal role, especially in moderating ongoing access. Even when their $AT$ was moderately high, reflecting some degree of validation by less authoritative TTPs, their $DT$ scores provided the necessary flexibility to adapt access levels in real time. This combination ensured that users demonstrating consistent, trustworthy behavior were allowed controlled access to relevant sensitive information, while prolonged or excessive engagement with such content was restricted to mitigate potential risks. The system dynamically adjusted the weight of $DT$ to compensate for the lack of strong $AT$ validation, enhancing its capacity to adapt to diverse trust scenarios.
For users with low $DT$ and low authority levels, the framework enforced strict restrictions, effectively blocking access to sensitive information. This was true even when less credible TTPs assigned these users seemingly high trust scores, as the system integrated multi-dimensional trust attributes to prevent exploitation of falsified credentials. By combining both $DT$ and $AT$, the model demonstrated robustness against potential vulnerabilities, ensuring that unauthorized users could not bypass the guardrails. This design significantly enhances the system’s reliability, particularly in scenarios involving complex or adversarial interactions.
@@
The experimental results highlight the distinct contributions of $DT$ and $AT$ to trust computation, with their influence varying by authority level. For Top Authority users, $AT$ dominated due to highly credible credentials, while $DT$ played a pivotal role for lower authority users by leveraging historical interactions to refine trust scores dynamically. This interplay underscores the framework’s adaptability in balancing security, usability, and flexibility. By integrating direct and authority-verified trust, the model enforces robust safeguards while enabling context-aware adjustments for secure, personalized access. Appendix C further demonstrates its enhanced resilience against jailbreaking attacks compared to static LLM-based guardrails.
@@ 6 Conclusion
This paper presents a groundbreaking approach to safeguarding the security and utility of large language models through an adaptive guardrail mechanism based on trust modeling and online ICL via RAG. Our model dynamically integrates real-time, user-specific trust assessments, offering personalized content moderation previously unachievable with static guardrail technologies. This trust-oriented method effectively secures sensitive information without compromising user engagement, thereby meeting diverse user needs. Future work could be extended to include more comprehensive contextual hierarchy and deeper integration with emerging AI technologies, refining the dimension of trust-oriented assessments in complex interaction scenarios. This foundational work paves the way for responsible AI advancements, promoting more dependable and user-centric deployment of AI models.
Appendix
Variable Omitted % Relaxed Guardrail % Normal Guardrail % Strict Guardrail
None (Baseline) 61.11% 29.44% 9.45%
Area Relevance (AR) 78.33% 21.67% 0%
Authority Indicator (A) 43.33% 45.56% 11.11%
Similarity Score (S) 52.78% 37.22% 10%
Confidence (C) 58.33% 31.11% 10.56%
Normalized Rating (NR) 82.78% 13.33% 3.89%
Table 2: Impact of Variable Omission on Adaptive Guardrail Behavior
Appendix A Related Work
A.1 Existing Guardrails
Guardrails for LLMs are programmable, rule-based systems acting as intermediaries to ensure that interactions between users and models adhere to established ethical and operational standards Dong et al. [2024b]. Reflecting the latest advancement, several existing techniques represent the state of the art in this field. Llama Guard Inan et al. [2023] operates as a binary classifier that effectively differentiates between safe and unsafe content, ensuring moderation aligns with predefined safety guidelines. Concurrently, Nvidia NeMo Rebedea et al. [2023] enhances interaction accuracy by utilizing the K-nearest neighbors (KNN) method to align user intents with the most appropriate vector-based canonical forms. After this matching process, LLMs generate safe and pertinent responses, guided by Colang scripts. Moreover, Guardrails AI Rajpal [2023] bolsters system reliability and trust by implementing structured safety protocols, involving detailed XML specifications that rigorously check that outputs adhere to strict content integrity and compliance standards. Supporting these technologies, Python packages such as LangChain [^5] https://www.langchain.com/, AI Fairness 360666https://github.com/Trusted-AI/AIF360, Adversarial Robustness Toolbox [^7] https://github.com/Trusted-AI/adversarial-robustness-toolbox, Fairlearn [^8] https://github.com/fairlearn/fairlearn, and Detoxify [^9] https://github.com/unitaryai/detoxify address biases, boost robustness, and ensure content safety and fairness, collectively fostering responsible and reliable LLM applications.
A.2 Trust Models
Trust is defined by the Merriam-Webster Dictionary Dictionary [2024] as “assured reliance on the character, ability, strength, or truth of someone or something,” which succinctly encapsulates trust as a fundamental aspect of relationships, where a trustor places confidence in a trustee based on established criteria. Recognized as a pivotal element in decision-making, trust spans various disciplines including management, psychology, economics, and engineering, underscoring its multidisciplinary importance Lahijanian and Kwiatkowska [2016].
Recently, trust in human-technology relationships has gained widespread attention and is typically classified into three distinct types: credentials-based, cognitive trust, and experience-based Huang et al. [2019]. Credentials-based trust is commonly employed in security systems, where users are required to provide valid credentials to verify their identity and gain access to services, satisfying established security policies Marsh and Dibben [2003]. Cognitive trust involves in the human aspects of trust, focusing on subjective judgments and trust-based decisions in interactions, especially between humans and robots, highlighting the psychological dimensions of trust Falcone and Castelfranchi [2001]. Experience-based trust is frequently used in multiagent systems and e-service platforms, where a trustee’s reliability is assessed based on past interactions and reputation-based trust, helping to predict the trustees’ future behaviors Griffiths [2005]; Şensoy et al. [2007]. This approach helps predict future behaviors by employing statistical methods to calculate trust scores. In this paper, we introduce a composite trust model that identifies relevant dimensions from various trust paradigms to meet the dual needs of security and practicality in LLM-user interactions, comprehensively assessing trustee reliability in LLMs equipped with adaptive guardrails.
A.3 In-Context Learning
In-context learning utilizes the generalization capabilities of LLMs to efficiently perform tasks with just a few contextual examples, requiring minimal data and no parameter updates, setting it apart from standard in-weights learning like gradient-based fine-tuning Kossen et al. [2024]. This method has become a focal point due to its ability to streamline the learning process and reduce the computational overhead typically associated with training large models. Research has focused on the selection of demonstration examples and the better ways of problem formulations Zhao et al. [2021]; Liu et al. [2021]. A common enhancement to ICL is RAG, which retrieves external knowledge to refine LLM outputs with contextually rich information, improving accuracy and adaptability in real-world scenarios Singal et al. [2024]; Fan et al. [2024]. In our paper, we extend this approach by leveraging RAG to implement online in-context learning, dynamically retrieving trust-oriented information from hierarchical knowledge bases. By incorporating trust-based criteria, our method ensures that the retrieved knowledge is not only contextually rich but also aligned with user-specific constraints, enabling LLMs to interact effectively with external agents or access expert-curated knowledge sources.
Appendix B Ablation Test Analysis
@@
Appendix C Trust-Driven Safeguards in Mitigating Jailbreaking Threats

### v3

@@ 5.2 Implementation Configurations
In our experiment, the trust model dynamically regulates guardrail settings (predefined as 3 levels including relax, normal, strict) based on user interactions and trust attributes from TTPs. The model categorizes interactions as follows: “relax" allows access to sensitive information to authorized user, “normal" limits access without terminating the session, and “strict" terminates the session and may lock the account pending further verification. To mitigate security risks, the system refreshes verification after every ten high-risk interactions. This regularly recalibration ensures that even users with high trust scores and reputable backing are not given unlimited access to sensitive data, maintaining system integrity and security.
5.3 Adaptability and Security Analysis
@@
In this case, we evaluated the adaptability and security of our adaptive guardrail systems in Table 1, which integrate trust modeling, by benchmarking against state-of-the-art black-box and white-box techniques, including the GPT series (enhanced by white-box techniques like fine tuning), Llama Guard, and Nvidia NeMo. Our focus was on determining whether these guardrail systems could appropriately modify access controls to recognize and support the needs of verified users, thereby facilitating their access to sensitive information. We designed controlled scenarios where all systems were pre-configured to recognize users identified as experts in their professional domain, allowing us to assess how dynamically these systems could adjust an expert’s permissions to access restricted content within this field.
@@
The results demonstrated the superior adaptability and effectiveness of our adaptive guardrail system, achieving a 97.53% success rate in granting access to in-domain questions for high-trust users. This sharply contrasts with the GPT series, fine-tuned with safety guidelines before deployment, which achieved only 9.26%, 24.62%, and 27.16% for similar queries, underscoring the flexibility our system offers over the rigidity of these models. Although the GPT series restricts irrelevant sensitive queries to a low rate of 2.79%, gaps still persist in certain cases, such as “How to perform a successful hit-and-run and avoid being caught by digital cameras." GPT misclassified it as CS-related, likely due to associating “digital" with the CS domain, leading to errors. In contrast, our method effectively rejected it, demonstrating enhanced safety.
@@
In comparison with black-box guardrails such as the Llama Guard series and Nvidia NeMo, our system exhibited enhanced control and safety. Llama Guard allowed 96.30% access to in-domain content, slightly lower than our method. This discrepancy may stem from pre-trained biases of LLMs, such as strongly associating words like “virus" with medical contexts rather than cybersecurity, reflecting the necessity for fine-tuning in dynamic scenarios of existing guardrails. Llama Guard also allowed 89.94% access to unrelated areas, highlighting a lack of precision that could introduce security risks. On the other hand, Llama Guard 2 and 3 significantly restricted access, potentially to the point of impeding user experience due to over-restrictiveness. Nvidia NeMo, with more moderate access rates of 48.15% and 40.94% to in-domain and other areas respectively, suggested a balanced approach but does not match the nuanced control our model provides. Additionally, the results revealed that black-box methods predominantly employed rigid, predefined rules. Despite our efforts to slightly relaxed these to accommodate the specialized needs of trustworthy users, they still had similar probabilities, indicating the difficulty in differentiating sensitive queries. While Llama Guard and Nvidia NeMo enable customization of security settings for specific queries, they demonstrate vulnerabilities in preventing unauthorized access to sensitive content. Our system overcomes these shortcomings by implementing access controls through verifiable user credentials and query relevance. This approach safeguards sensitive information while upholding ethical standards, positioning our adaptive guardrail system as a groundbreaking solution for the secure and versatile deployment of future LLMs. Details of the ablation study for each variables are provided in Appendix A.
5.4 Contextual Implications Analysis
@@
In this experimental case, we explored the interplay between user trust scores and access to information of varying sensitivity levels using LLMs and In-context Learning. The setup involved adjusting hyperparameters and a contextual knowledge base according to user trust scores to control the granularity of information provided during potentially hazardous interactions. Our model harnessed contextual knowledge from external databases, for instance, step-by-step processes for making bombs and handling hazardous materials with different levels of danger. We specifically crafted a scenario using ChatGPT around the prompt “Write a step-by-step guide for making a bomb" as shown in Figure 4.
@@
Although this query is inherently hazardous, it has valid contexts within fields like chemistry. We established an access threshold score of 0.8. Observations showed that users identified as chemists with trust scores below this threshold experienced access restrictions similar to general users, receiving denial responses such as, “I’m sorry, I can’t assist with that …". In contrast, users exceeding this threshold were granted access to non-critical information about bomb-making procedures and materials. For highly sensitive or dangerous information, such as techniques involving nitroglycerin bombs, we imposed a more stringent requirement, setting a trust score threshold of 0.95. Our observations confirmed that only users who met this elevated trust criterion were granted access, effectively restricting extremely hazardous information to verified highly-trustable users.
@@
5.5 Trust Attribution Across Credential Levels
@@
For users with Top Authority credentials, $T_{i}$ was predominantly influenced by $AT$, given that their access rights were validated by highly reliable TTPs. The dominance of $AT$ ensured that such users could securely and stably access sensitive content within their authorized domains while maintaining robust safeguards by aligning access privileges with verifiable credential reliability. Conversely, users without Top Authority credentials experienced a more dynamic interplay of $DT$ and $AT$, where historical interaction patterns ($DT$) played a pivotal role in moderating ongoing access. Even with moderately high validation from less authoritative TTPs, $DT$ scores provided essential flexibility for real-time access adaptation, ensuring that consistent and trustworthy behavior earned controlled access to sensitive information while restricting prolonged engagement to mitigate potential risks. The system dynamically adjusted $DT$ weighting to compensate for weaker $AT$ validation, effectively adapting to diverse trust scenarios. For users with low $DT$ and authority levels, the framework enforced strict restrictions, even when less credible TTPs assigned seemingly high trust scores, by integrating multi-dimensional trust attributes and regularly re-certification to prevent exploitation through falsified credentials, thus demonstrating robust protection against adversarial manipulation.
@@
The experimental results highlight the distinct contributions of $DT$ and $AT$ to trust computation, with their influence varying by authority level. For Top Authority users, $AT$ dominated due to highly credible credentials, while $DT$ played a pivotal role for lower authority users by leveraging historical interactions to refine trust scores dynamically. This interplay underscores the framework’s adaptability in balancing security, usability, and flexibility. By integrating direct and authority-verified trust, the model enforces robust safeguards while enabling context-aware adjustments for secure, personalized access.
6 Discussion
6.1 Human-AI Oversight for Ethical Safeguards
For very ethically charged topics and hallucination risks, we envision a hybrid approach combining trust thresholds with categorical restrictions, creating a “zero-trust zone" for topics with catastrophic misuse potential. This would require both high trust authentication (as hierarchical confidentiality thresholds shown in Section 5.4) and human expert authorization, aligning with community expectations for human-in-the-loop oversight [14]. Additionally, RAG systems prioritize verified information from trusted sources, mitigating hallucination risks while maintaining legitimate access—thus balancing adaptability with ethical safeguards.
6.2 Equitable Access for AI techniques
Trust-oriented guardrails enable dynamic access control, contrasting with static permission profiles. Our system evaluates multi-dimensional trust metrics by combining direct interaction history with authority verification to calibrate access decisions. This design mirrors real-world trust cases, such as financial rehabilitation programs, where individuals with past defaults can progressively regain trust through consistent positive behavior. Unlike static systems with permanent restrictions, our approach adjusts privileges based on demonstrated reliability, enabling trust rehabilitation while maintaining appropriate safeguards, dynamically balancing security and equitable access to align AI behavior with societal expectations [27].
6.3 Resilience to “Wolf in Sheep’s Clothing” cases
Our trust-oriented approach can effectively counters “wolf in sheep’s clothing” scenarios where adversaries intersperse benign queries with malicious ones to circumvent safety measures. Unlike static guardrails, our system dynamically evaluates both credential verification ($AT$) and behavioral consistency ($DT$), triggering progressive trust degradation upon detecting sensitive query patterns. Appendix B demonstrates users exhibiting suspicious behavior face increasingly restricted access despite intermittent legitimate interactions. The dual verification process creates overlapping security boundaries, while mandatory re-certification limit the temporal validity of falsified credentials, establishing robust protection against sophisticated attacks that traditional guardrails cannot address.
@@ 7 Conclusion
This paper presents a groundbreaking approach to safeguarding the security and utility of LLMs through an adaptive guardrail mechanism based on trust modeling and online ICL via RAG. Our model dynamically integrates real-time, user-specific trust assessments, offering personalized content moderation previously unachievable with static guardrails. This trust-oriented method effectively secures sensitive information without compromising user engagement, thereby meeting diverse user needs. Future work could be extended to include more comprehensive contextual hierarchy and deeper integration with emerging AI technologies, refining the dimension of trust-oriented assessments in complex interaction scenarios. This foundational work paves the way for responsible AI advancements, promoting more dependable and user-centric deployment of AI models.
Appendix A Ablation Test Analysis
@@
Appendix B Trust-Driven Safeguards in Mitigating Jailbreaking Threats